OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[2\sqrt{50}-3\sqrt{8}\]

jimthompson5910 (jim_thompson5910):

\[2\sqrt{50}-3\sqrt{8}\] \[2\sqrt{25*2}-3\sqrt{4*2}\] \[2\sqrt{25}*\sqrt{2}-3\sqrt{4}*\sqrt{2}\] \[2*5*\sqrt{2}-3*2*\sqrt{2}\] \[10*\sqrt{2}-6*\sqrt{2}\] \[4\sqrt{2}\]